b. XYZ (Pty) Ltd commenced business on 1st January 2014. On that day, XYZ (Pty)
Ltd bought equipment on credit from CCC equipment (Pty) Ltd at P85 000. The
equipment will be used for 5 years and the estimated value at the end of the 5
years is P25, 000. Calculate the depreciation for the equipment at 10% for the first
three years using the reducing balance method. (5 marks)
Answer #1

Reducing balance method of depreciation

(Amount in P)

Year Opening WDV Depreciation Closing WDV
1                  85,000                     6,000                79,000
2                  79,000                     5,400                73,600
3                  73,600                     4,860                68,740
Rate of Depreciation 10%
Useful Life 5
Residual Value 25000
WDV Written Down Value

--

Formula:

D Opening WDV 85000 =E6 =E7 Depreciation =(C6-$C$12) * $C$10 =(C7-$C$12)*$C$10 =(C8-$C$12) * $C$10 Closing WDV =C6-D6 =C7-D7

--

Hope you Understood.
If you have any doubt please leave a comment.

Thank you.